### v1

@@ Abstract
The performance of deep learning models in remote sensing (RS) strongly depends on the availability of high-quality labeled data. However, collecting large-scale annotations is costly and time-consuming, while vast amounts of unlabeled imagery remain underutilized. To address this challenge, we propose a Hierarchical Semi-Supervised Active Learning (HSSAL) framework that integrates semi-supervised learning (SSL) and a novel hierarchical active learning (HAL) in a closed iterative loop. In each iteration, SSL refines the model using both labeled data through supervised learning and unlabeled data via weak-to-strong self-training, improving feature representation and uncertainty estimation. Guided by the refined representations and uncertainty cues of unlabeled samples, HAL then conducts sample querying through a progressive clustering strategy, selecting the most informative instances that jointly satisfy the criteria of scalability, diversity, and uncertainty. This hierarchical process ensures both efficiency and representativeness in sample selection. Extensive experiments on three benchmark RS scene classification datasets, including UCM, AID, and NWPU-RESISC45, demonstrate that HSSAL consistently outperforms SSL- or AL-only baselines. Remarkably, with only 8%, 4%, and 2% labeled training data on UCM, AID, and NWPU-RESISC45, respectively, HSSAL achieves over 95% of fully-supervised accuracy, highlighting its superior label efficiency through informativeness exploitation of unlabeled data. Our code will be released at https://github.com/zhu-xlab/RS-SSAL.
keywords:
[…]
1 Introduction
@@
The performance of deep learning models in remote sensing (RS) is heavily limited by the scarcity of high-quality labeled samples, which are costly and time-consuming to collect, often requiring expert knowledge. In contrast, Earth observation satellites continuously generate vast amounts of unlabeled imagery. Thus, a central challenge is to develop learning paradigms that effectively mitigate label scarcity by extracting greater utility from the abundant unlabeled data. Addressing this challenge is critical for enabling efficient and scalable applications in RS tasks such as scene classification (cheng2017remote; wang2020looking; huang2023semi), land-use classification (karra2021global; wang2023review), object detection (li2017rotation; zhang2023efficient; zhang2024structured), and semantic segmentation (diakogiannis2020resunet; hua2021semantic; huang2024decouple).
@@
To this end, semi-supervised learning (SSL(oliver2018realistic; sohn2020fixmatch; zhang2021flexmatch; wang2023freematch) and active learning(AL(settles2009active; yoo2019learning; bald; coreset; badge) offer complementary yet independently limited strategies. SSL enhances model performance by utilizing unlabeled data for model self-training, but its effectiveness depends significantly on the quality and representativeness of the initial labeled set, which is often randomly constructed and may lead to suboptimal learning. AL, on the other hand, mitigates annotation scarcity by iteratively selecting the most informative samples based on criteria such as diversity or representativeness for human labeling. However, by focusing only on a small subset of instances in each cycle, conventional AL fails to utilize the vast majority of unlabeled data, which remain unused in model training.
@@
The integration of these two strategies gives rise to semi-supervised active learning (SSAL(huang2021semi; guo2021semi; gao2020consistency), a unified paradigm in which SSL and AL complement and reinforce each other in a synergistic loop. In SSAL, the SSL component leverages both labeled data via supervised learning and unlabeled data through pseudo-labeling to gradually enhance the model’s representation quality and uncertainty estimation. Following the SSL phase, the AL module identifies and queries the most informative unlabeled samples for manual annotation, thereby expanding the labeled set in a targeted and cost-effective manner. Importantly, the two modules interact bidirectionally: a more discriminative SSL-trained model provides more discriminative representation and more reliable uncertainty cues to guide AL toward better query decisions, while the carefully selected samples from AL, in turn, supply higher-value labels that further strengthen SSL training in the next round. This mutual reinforcement enables SSAL to continuously improve both model performance and querying effectiveness. Despite its strong potential to reduce annotation costs and improve data efficiency, SSAL remains largely underexplored in the RS domain.
@@ 2 Related Work
2.1 Semi-Supervised Learning
2.1.1 General Semi-Supervised Learning
@@
SSL has gained increasing attention in the deep learning era, as it effectively alleviates the annotation burden when labeled data are scarce. Early deep learning–based SSL methods leveraged adversarial learning, particularly GANs (souly2017semi; hung2018adversarial; mittal2019semi), to align feature distributions between labeled and unlabeled data via a discriminator that minimizes prediction entropy and domain discrepancy. While effective to some extent, these adversarial approaches suffer from poor interpretability and unstable optimization.
@@
Subsequently, consistency-based methods emerged as a more stable alternative, enforcing prediction invariance under input or feature perturbations (ouali2020semi; liu2022perturbed; sohn2020fixmatch; chen2021semi). By encouraging consistent outputs for augmented views of the same sample, they enhance representation robustness and generalization. However, consistency-based methods often lack effective pseudo-label refinement, limiting iterative improvement through high-confidence predictions.
@@
Recently, self-training–based methods have become dominant due to their strong performance and simplicity. These approaches generate pseudo-labels for unlabeled data and jointly optimize supervised and self-training losses. Representative methods include FixMatch (sohn2020fixmatch), which filters pseudo-labels with a high-confidence threshold; FlexMatch (zhang2021flexmatch), which adjusts class-wise thresholds dynamically; FreeMatch (wang2023freematch), which uses adaptive confidence estimation to reduce overfitting; SoftMatch (chen2023softmatch), which applies soft weighting for uncertain samples; and CGMatch (cheng2025cgmatch), which partitions unlabeled samples into easy, ambiguous, and hard subsets for selective regularization. These methods collectively highlight the importance of confidence- and uncertainty-aware learning in SSL.
2.1.2 Semi-Supervised Learning in Remote Sensing
@@
SSL has also been widely applied in RS to alleviate the annotation burden on massive unlabeled imagery. For RS scene classification, Miao et al. (miao2022semi) proposed a semi-supervised representation consistency Siamese network combining Involution-GAN feature extraction with a Siamese consistency loss. Ding et al. (ding2024uncertainty) introduced uncertainty-aware contrastive learning to handle reliable and unreliable samples differently, improving multimodal classification. Liu et al. (liu2023multi) proposed a multi-level label-aware framework in unlabeled data for better pseudo-labeling and feature discrimination. Huang et al. (huang2023semi) developed a semi-supervised domain adaptation method for cross-domain RS scene classification, aligning source and target domains with both unsupervised and class-level supervised strategies. These advances improve RS-SSL by addressing pseudo-label noise, domain variability, and model uncertainty.
@@
Beyond scene classification, SSL has been explored in change detection (bandara2022revisiting; zhang2023joint; shen2024learning), road/building extraction (huang2023adaptmatch; huang2024semi), and RS semantic segmentation (sun2020bas; zhang2022semi; chen2022semi; huang2024decouple; sun2025rsprotosemiseg; wang2025semi), highlighting RS-specific challenges such as high intra-class diversity and inter-class similarity caused by heterogeneous ground and observation conditions.
2.2 Active and Semi-Supervised Active Learning
@@
AL aims to enhance model performance under limited supervision by selectively querying the most informative samples for annotation, thereby maximizing the efficiency of labeling resources. Traditional AL methods generally rely on uncertainty-based or diversity-based sampling strategies. Uncertainty-based approaches (lewis1995sequential; settles2009active; joshi2009multi) select samples on which the model is least confident, often measured via prediction entropy or margin sampling. Diversity-based methods (coreset; badge; agarwal2020contextual) seek to ensure representativeness by covering diverse regions of the feature space, mitigating redundancy among queried samples. Recent deep AL frameworks further integrate these criteria with modern neural architectures: for instance, CoreSet (coreset) formulates AL as a core-set selection problem via k-center clustering; BADGE (badge) combines uncertainty and diversity by selecting gradient-embedding-based samples; and VAAL (sinha2019variational) employs adversarial variational autoencoders to discriminate labeled and unlabeled distributions for sample selection. Uncertainty-aware extension (yoo2019learning) introduces deep uncertainty estimation and achieves improved performance with fewer labeled samples.
@@
In the RS domain, AL has been extensively adopted to alleviate the high annotation cost associated with large-scale satellite and aerial imagery. Early studies (tuia2009active; persello2014active; stumpf2013active; demir2010batch) primarily focused on uncertainty- or region-based querying for hyperspectral image (HSI) and land-cover classification, effectively identifying the most informative samples under limited labeling budgets. To further enhance the reliability of sample selection, subsequent works incorporated spatial–spectral context (patra2017spectral; tuia2011survey) and multi-view representations (xu2021multiview) to better handle the heterogeneity of RS data. With the advent of deep learning, AL has evolved toward feature-aware and structure-guided strategies. For instance, Liu et al. (lei2021active) proposed a deep AL framework for HSI classification that integrates spectral–spatial uncertainty with manifold regularization, achieving superior generalization under scarce supervision. These advances collectively demonstrate that AL can substantially reduce labeling costs across diverse RS applications, including HSI analysis (liu2018superpixel; cai2021phase), scene classification (stumpf2013active; demir2010batch), and semantic segmentation (lenczner2022dial; desai2022active).
@@
To exploit the remaining unlabeled data more effectively, SSAL unifies AL’s selective querying with SSL’s ability to reduce the uncertainty of unlabeled samples, thereby achieving superior label efficiency. Representative examples in the general domain include ConsistencyAL (gao2020consistency), which enforces prediction consistency between labeled and unlabeled data to guide query selection; SemiAL (guo2021semi), which alternates between pseudo-label–based self-training and uncertainty-driven annotation; and S3AL (huang2021semi), which jointly optimizes sampling and model learning in a single semi-supervised objective. Compared with prior SSAL frameworks such as ConsistencyAL, SemiAL, and S3AL that mainly rely on flat uncertainty- or consistency-based querying, HSSAL employs a hierarchical selection mechanism that jointly enforces scalability, diversity, and uncertainty, which is particularly effective for RS scene classification with highly non-convex and imbalanced class manifolds.
3 Hierarchical Semi-Supervised Active Learning
3.1 Overview of HSSAL with Dataset Notations
@@
3.2 Semi-Supervised Learning
The SSL component in HSSAL is built upon a weak-to-strong self-training mechanism originating from FixMatch (sohn2020fixmatch), which enables the model to effectively leverage the large pool of unlabeled data by generating and refining pseudo-labels for self-supervised training. This mechanism represents one of the most classical SSL paradigms, allowing the model to jointly learn from a small set of labeled samples and a large set of pseudo-labeled samples produced from its own predictions under weak augmentations.
@@ 3.2 Semi-Supervised Learning
3.3 Hierarchical Active Learning
@@
3.3.1 Step I: Feature Extraction and Uncertainty Estimation
We decompose the EMA model into an encoder–classifier architecture with parameters $\hat{\theta}=(\hat{\theta}_{\text{enc}},W)$, where $\hat{\theta}_{\text{enc}}$ denotes the DINOv2-Small encoder and $W$ is the final linear classifier. HAL leverages the encoder features for clustering and the classifier gradients for uncertainty estimation, following the principle of BADGE (badge) but without performing gradient-space clustering. Building upon the well-trained EMA model obtained from the preceding SSL phase, HAL characterizes the entire unlabeled set $\mathcal{U}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ to derive two complementary representations: a global feature matrix $\mathbf{F}$ that captures the semantic distribution, and an uncertainty vector $\mathbf{u}$ that quantifies sample informativeness. These two components jointly form the basis for the subsequent mini-batch partitioning and uncertainty-aware querying process.
@@ 3.3.1 Step I: Feature Extraction and Uncertainty Estimation
Unlike conventional entropy-based or margin-based methods, we adopt the gradient norm of the classifier as the uncertainty metric, following (badge). For each sample $x_{i}$, the uncertainty score $u_{i}$ is computed as:
@@
3.3.2 Step II: Mini-batch Partition for Scalable Selection
@@
3.3.3 Step III: Spectral Clustering and Uncertainty-aware Sampling
@@
3.3.4 Step IV: Sample Annotation and Dataset Update
@@ 4 Experiments
4.1 Datasets, metrics, and experimental settings
Datasets. To comprehensively evaluate the effectiveness of the proposed HSSAL framework for RS scene classification, extensive experiments were conducted on three widely used benchmarks: (1) UCM (UCM) contains 2,100 aerial images of size $256\times 256$ pixels across 21 land-cover categories such as residential, forest, and airport scenes, with a spatial resolution of approximately 0.3 m; (2) AID (AID) comprises 10,000 images of size $600\times 600$ pixels collected from Google Earth, spanning 30 scene types with high intra-class variation and inter-class similarity, at a spatial resolution ranging from 0.5 m to 8 m; and (3) NWPU-RESISC45 (RESISC45) consists of 31,500 images of size $256\times 256$ pixels over 45 scene classes, representing one of the largest and most diverse RS scene classification benchmarks to date, with spatial resolutions varying between 0.2 m and 30 m. All datasets provide RGB imagery under diverse spatial resolutions and geographic distributions, allowing for a comprehensive evaluation of the model’s generalization capability.

### v2

@@ Abstract
The performance of deep learning models in remote sensing (RS) strongly depends on the availability of high-quality labeled data. However, collecting large-scale annotations is costly and time-consuming, while vast amounts of unlabeled imagery remain underutilized. To address this challenge, we propose a Hierarchical Semi-Supervised Active Learning (HSSAL) framework that integrates semi-supervised learning (SSL) and a novel hierarchical active learning (HAL) in a closed iterative loop. In each iteration, SSL refines the model using both labeled data through supervised learning and unlabeled data via weak-to-strong self-training, improving feature representation and uncertainty estimation. Guided by the refined representations and uncertainty cues of unlabeled samples, HAL then conducts sample querying through a progressive clustering strategy, selecting the most informative instances that jointly satisfy the criteria of scalability, diversity, and uncertainty. This hierarchical process ensures both efficiency and representativeness in sample selection. Extensive experiments on three benchmark RS scene classification datasets, including UCM, AID, and NWPU-RESISC45, demonstrate that HSSAL consistently outperforms SSL- or AL-only baselines. Remarkably, with only 8%, 4%, and 2% labeled training data on UCM, AID, and NWPU-RESISC45, respectively, HSSAL achieves over 95% of fully-supervised accuracy, highlighting its superior label efficiency through informativeness exploitation of unlabeled data. Our code will be publicly available.
I Introduction
@@
The performance of deep learning models in remote sensing (RS) is heavily limited by the scarcity of high-quality labeled samples, which are costly and time-consuming to collect, often requiring expert knowledge. In contrast, Earth observation satellites continuously generate vast amounts of unlabeled imagery. Thus, a central challenge is to develop learning paradigms that effectively mitigate label scarcity by extracting greater utility from the abundant unlabeled data. Addressing this challenge is critical for enabling efficient and scalable applications in RS tasks such as scene classification[1, 2, 3], land-use classification[4, 5], object detection[6, 7, 8], and semantic segmentation[9, 10, 11].
@@
To this end, semi-supervised learning (SSL)[12, 13, 14, 15] and active learning(AL)[16, 17, 18, 19, 20] offer complementary yet independently limited strategies. SSL enhances model performance by utilizing unlabeled data for model self-training, but its effectiveness depends significantly on the quality and representativeness of the initial labeled set, which is often randomly constructed and may lead to suboptimal learning. AL, on the other hand, mitigates annotation scarcity by iteratively selecting the most informative samples based on criteria such as diversity or representativeness for human labeling. However, by focusing only on a small subset of instances in each cycle, conventional AL fails to utilize the vast majority of unlabeled data, which remain unused in model training.
@@
The integration of these two strategies gives rise to semi-supervised active learning (SSAL)[21, 22, 23], a unified paradigm in which SSL and AL complement and reinforce each other in a synergistic loop. In SSAL, the SSL component leverages both labeled data via supervised learning and unlabeled data through pseudo-labeling to gradually enhance the model’s representation quality and uncertainty estimation. Following the SSL phase, the AL module identifies and queries the most informative unlabeled samples for manual annotation, thereby expanding the labeled set in a targeted and cost-effective manner. Importantly, the two modules interact bidirectionally: a more discriminative SSL-trained model provides more discriminative representation and more reliable uncertainty cues to guide AL toward better query decisions, while the carefully selected samples from AL, in turn, supply higher-value labels that further strengthen SSL training in the next round. This mutual reinforcement enables SSAL to continuously improve both model performance and querying effectiveness. Despite its strong potential to reduce annotation costs and improve data efficiency, SSAL remains largely underexplored in the RS domain.
@@ II Related Work
II-A Semi-Supervised Learning
II-A1 General Semi-Supervised Learning
@@
SSL has gained increasing attention in the deep learning era, as it effectively alleviates the annotation burden when labeled data are scarce. Early deep learning–based SSL methods leveraged adversarial learning, particularly GANs[24, 25, 26], to align feature distributions between labeled and unlabeled data via a discriminator that minimizes prediction entropy and domain discrepancy. While effective to some extent, these adversarial approaches suffer from poor interpretability and unstable optimization.
@@
Subsequently, consistency-based methods emerged as a more stable alternative, enforcing prediction invariance under input or feature perturbations[27, 28, 13, 29]. By encouraging consistent outputs for augmented views of the same sample, they enhance representation robustness and generalization. However, consistency-based methods often lack effective pseudo-label refinement, limiting iterative improvement through high-confidence predictions.
@@
Recently, self-training–based methods have become dominant due to their strong performance and simplicity. These approaches generate pseudo-labels for unlabeled data and jointly optimize supervised and self-training losses. Representative methods include FixMatch[13], which filters pseudo-labels with a high-confidence threshold; FlexMatch[14], which adjusts class-wise thresholds dynamically; FreeMatch[15], which uses adaptive confidence estimation to reduce overfitting; SoftMatch[30], which applies soft weighting for uncertain samples; and CGMatch[31], which partitions unlabeled samples into easy, ambiguous, and hard subsets for selective regularization. These methods collectively highlight the importance of confidence- and uncertainty-aware learning in SSL.
II-A2 Semi-Supervised Learning in Remote Sensing
@@
SSL has also been widely applied in RS to alleviate the annotation burden on massive unlabeled imagery. For RS scene classification, Miao et al.[32] proposed a semi-supervised representation consistency Siamese network combining Involution-GAN feature extraction with a Siamese consistency loss. Ding et al.[33] introduced uncertainty-aware contrastive learning to handle reliable and unreliable samples differently, improving multimodal classification. Liu et al.[34] proposed a multi-level label-aware framework in unlabeled data for better pseudo-labeling and feature discrimination. Huang et al.[3] developed a semi-supervised domain adaptation method for cross-domain RS scene classification, aligning source and target domains with both unsupervised and class-level supervised strategies. These advances improve RS-SSL by addressing pseudo-label noise, domain variability, and model uncertainty.
@@
Beyond scene classification, SSL has been explored in change detection[35, 36, 37], road/building extraction[38, 39], and RS semantic segmentation[40, 41, 42, 11, 43, 44], highlighting RS-specific challenges such as high intra-class diversity and inter-class similarity caused by heterogeneous ground and observation conditions.
II-B Active and Semi-Supervised Active Learning
@@
AL aims to enhance model performance under limited supervision by selectively querying the most informative samples for annotation, thereby maximizing the efficiency of labeling resources. Traditional AL methods generally rely on uncertainty-based or diversity-based sampling strategies. Uncertainty-based approaches[45, 16, 46] select samples on which the model is least confident, often measured via prediction entropy or margin sampling. Diversity-based methods[19, 20, 47] seek to ensure representativeness by covering diverse regions of the feature space, mitigating redundancy among queried samples. Recent deep AL frameworks further integrate these criteria with modern neural architectures: for instance, CoreSet[19] formulates AL as a core-set selection problem via k-center clustering; BADGE[20] combines uncertainty and diversity by selecting gradient-embedding-based samples; and VAAL[48] employs adversarial variational autoencoders to discriminate labeled and unlabeled distributions for sample selection. Uncertainty-aware extension[17] introduces deep uncertainty estimation and achieves improved performance with fewer labeled samples.
@@
In the RS domain, AL has been extensively adopted to alleviate the high annotation cost associated with large-scale satellite and aerial imagery. Early studies[49, 50, 51, 52] primarily focused on uncertainty- or region-based querying for hyperspectral image (HSI) and land-cover classification, effectively identifying the most informative samples under limited labeling budgets. To further enhance the reliability of sample selection, subsequent works incorporated spatial–spectral context[53, 54] and multi-view representations[55] to better handle the heterogeneity of RS data. With the advent of deep learning, AL has evolved toward feature-aware and structure-guided strategies. For instance, Liu et al.[56] proposed a deep AL framework for HSI classification that integrates spectral–spatial uncertainty with manifold regularization, achieving superior generalization under scarce supervision. These advances collectively demonstrate that AL can substantially reduce labeling costs across diverse RS applications, including HSI analysis[57, 58], scene classification[51, 52], and semantic segmentation[59, 60].
@@
To exploit the remaining unlabeled data more effectively, SSAL unifies AL’s selective querying with SSL’s ability to reduce the uncertainty of unlabeled samples, thereby achieving superior label efficiency. Representative examples in the general domain include ConsistencyAL[23], which enforces prediction consistency between labeled and unlabeled data to guide query selection; SemiAL[22], which alternates between pseudo-label–based self-training and uncertainty-driven annotation; and S3AL[21], which jointly optimizes sampling and model learning in a single semi-supervised objective. Compared with prior SSAL frameworks such as ConsistencyAL, SemiAL, and S3AL that mainly rely on flat uncertainty- or consistency-based querying, HSSAL employs a hierarchical selection mechanism that jointly enforces scalability, diversity, and uncertainty, which is particularly effective for RS scene classification with highly non-convex and imbalanced class manifolds.
III Hierarchical Semi-Supervised Active Learning
III-A Overview of HSSAL with Dataset Notations
@@
III-B Semi-Supervised Learning
The SSL component in HSSAL is built upon a weak-to-strong self-training mechanism originating from FixMatch[13], which enables the model to effectively leverage the large pool of unlabeled data by generating and refining pseudo-labels for self-supervised training. This mechanism represents one of the most classical SSL paradigms, allowing the model to jointly learn from a small set of labeled samples and a large set of pseudo-labeled samples produced from its own predictions under weak augmentations.
@@ III-B Semi-Supervised Learning
III-C Hierarchical Active Learning
@@
III-C1 Step I: Feature Extraction and Uncertainty Estimation
We decompose the EMA model into an encoder–classifier architecture with parameters $\hat{\theta}=(\hat{\theta}_{\text{enc}},W)$, where $\hat{\theta}_{\text{enc}}$ denotes the DINOv2-Small encoder and $W$ is the final linear classifier. HAL leverages the encoder features for clustering and the classifier gradients for uncertainty estimation, following the principle of BADGE[20] but without performing gradient-space clustering. Building upon the well-trained EMA model obtained from the preceding SSL phase, HAL characterizes the entire unlabeled set $\mathcal{U}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ to derive two complementary representations: a global feature matrix $\mathbf{F}$ that captures the semantic distribution, and an uncertainty vector $\mathbf{u}$ that quantifies sample informativeness. These two components jointly form the basis for the subsequent mini-batch partitioning and uncertainty-aware querying process.
@@ III-C1 Step I: Feature Extraction and Uncertainty Estimation
Unlike conventional entropy-based or margin-based methods, we adopt the gradient norm of the classifier as the uncertainty metric, following[20]. For each sample $x_{i}$, the uncertainty score $u_{i}$ is computed as:
@@
III-C2 Step II: Mini-batch Partition for Scalable Selection
@@
III-C3 Step III: Spectral Clustering and Uncertainty-aware Sampling
@@
III-C4 Step IV: Sample Annotation and Dataset Update
@@ IV Experiments
IV-A Datasets, metrics, and experimental settings
Datasets. To comprehensively evaluate the effectiveness of the proposed HSSAL framework for RS scene classification, extensive experiments were conducted on three widely used benchmarks: (1) UCM[61] contains 2,100 aerial images of size $256\times 256$ pixels across 21 land-cover categories such as residential, forest, and airport scenes, with a spatial resolution of approximately 0.3 m; (2) AID[62] comprises 10,000 images of size $600\times 600$ pixels collected from Google Earth, spanning 30 scene types with high intra-class variation and inter-class similarity, at a spatial resolution ranging from 0.5 m to 8 m; and (3) NWPU-RESISC45[63] consists of 31,500 images of size $256\times 256$ pixels over 45 scene classes, representing one of the largest and most diverse RS scene classification benchmarks to date, with spatial resolutions varying between 0.2 m and 30 m. All datasets provide RGB imagery under diverse spatial resolutions and geographic distributions, allowing for a comprehensive evaluation of the model’s generalization capability.